OpenStudy (anonymous):

(9/x^2-6x+9)+(2/x^2-9) how do I add this?

OpenStudy (anonymous):

\[ \frac{9}{x^2-6 x+9}+\frac{2}{x^2-9}=\frac{2}{(x-3) (x+3)}+\frac{9}{(x-3)^2}=\frac{11 x+21}{(x-3)^2 (x+3)} \]

OpenStudy (anonymous):

gracias

OpenStudy (anonymous):

w